LHC 2010 - latest news

Week 39: Machine coordinators: Mike Lamont, Jan Uythoven, Malika Meddahi

Main aims of the week:


Tuesday afternoon

Injection optimization/16 bunch injection

Wednesday morning

Transverse damper studies
Feedback upgrade & test ramp

Thursday morning

Abort gap cleaning setup


van der Meer scans


Physics 152b


Physics 152b


LHC coordination vistar



Link to report from morning meeting

Monday 4th October

  • Preparing for the next intensity step

Slides from 8:30 meeting

Sunday 3rd October

  • Filling for 152b physics fill as soon as SPS ready with beam
  • Stable beams 152b

Slides from 9:00 meeting

Saturday 2nd October

  • Physics 152 bunches got dumped by QPS after 2h of stable beams
  • All pm shift : repair of QPS problem (linked to computer failure)
  • Night: recovering – SPS main PC failure

Slides from 9:00 meeting

Friday 1st October

  • Van der Meer scans
  • Physics 152 bunches – dumped by UFO after 2h10 of stable beams

Slides from 8:30 meeting

Thursday 30th September

  • Abort Gap cleaning studies
  • Injection vs Vacuum problems

Slides from 8:30 meeting

Wednesday 29th September

  • Recover QPS, PC issues
  • Injection setup 16 bunches
  • Physics 152 bunches

Slides from 8:30 meeting

Tuesday 28th September

  • 02:30 Fill 1373 - 104 bunches - initial luminosity 3.5e31
  • Morning: TOTEM RPs in (Stable beams)
  • 11:15 Beams dumped
  • Afternoon: injection setup (16 bunches per SPS extraction) curtailed by PS problems
  • Attempt for physics 152 bunches: UFO, QPS and PC prevented

Slides from 8:30 meeting

Monday 27th September

  • 09:30 End of fill measurements
  • 11:00 access for tune monitor (BBQ) and quench heater power supply repair
  • Afternoon: RF tests & recovery from access
  • 17:00 beam back
  • Lost sectors 67 & 78 - QPS glitch

Slides from 08:30 meeting